OpenStudy (kitkat77):

Evaluate. 90 • 4 – 14 ÷ 2 + 102 ÷ 5 • 2

OpenStudy (anonymous):

Just seperate them by parentheses, that makes it much easier. When seperating, keep pairs of multiplication and division seperate. Example: \[\large \sf (90 \times 4) - (\frac{14}{2}) + (\frac{102}{5} \times 2)\]

OpenStudy (kitkat77):

okay thank you.